A population consists of the 5 measurements 2, 6, 8, 0 and 1 and samples of size n = 2 . What is the variance of the population? 2 DECIMAL PLACES 



1
Expert's answer
2021-05-07T08:59:01-0400

Variance= ó2/n (N-n/N-1)

Ó2= summation X2/N -(summation X/N)2

X: 2 6 8 0 1 =17

X2:4 36 64 0 1=105

Ó2 =(105/5) - (17/5)

21-11.56=9.44

Var= 9.44/2(3/4)

= 3.54